HETAV PATEL

Grade 11,

The perimeter of a certain sector of a circle is equal to the length of the arc of a semi-circle having the same radius, the angle of the sector is A. 65o 24' B. 64o 24' C. 63o 24' D. 62o 24

The perimeter of a certain sector of a circle is equal to the length of the arc of a semi-circle having the same radius, the angle of the sector is
A. 65o 24'
B. 64o 24'
C. 63o 24'
D. 62o 24

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
perimeter of sector = 2\pirA/360 + 2r
Here, A = angle of sector
And, length of arc of semi-circle of same radius = \pir
Thus,
\pirA/180 + 2r = \pir
On solving,
A = 180(1 - 2/\pi) taking \pi= 22/7
We get, A = 720/11
Hence A is the correct answer
Thanks and Regards
shaik Aasif
askIITians faculty